# Continuous time asymptotic representations for adaptive experiments

### v1

@@ Abstract.
This article develops a continuous-time asymptotic framework for analyzing adaptive experiments—settings in which data collection and treatment assignment evolve dynamically in response to incoming information. A key challenge in analyzing fully adaptive experiments, where the assignment policy is updated after each observation, is that the sequence of policy rules often lack a well-defined asymptotic limit. To address this, we focus instead on the empirical allocation process, which captures the fraction of observations assigned to each treatment over time. We show that, under general conditions, any adaptive experiment and its associated empirical allocation process can be approximated by a limit experiment defined by Gaussian diffusions with unknown drifts and a corresponding continuous-time allocation process. This limit representation facilitates the analysis of optimal decision rules by reducing the dimensionality of the state-space and leveraging the tractability of Gaussian diffusions. We apply the framework to derive optimal estimators, analyze in-sample regret for adaptive experiments, and construct e-processes for anytime-valid inference. Notably, we introduce the first definition of any-time and any-experiment valid inference for multi-treatment settings.
This version: January 2, 2026
@@ 1. Introduction
Adaptive experiments are experiments where data is collected and analyzed continuously, allowing for adjustments or decisions to be made on an ongoing basis. Originating from early work by wald1947sequential, arrow1949bayes, and others, adaptive experimentation has evolved to encompass a wide range of strategies, including bandit experiments, A/B testing, costly sampling, and best-arm identification. These strategies are now widely used across various fields. For instance, tech companies frequently employ bandit algorithms and A/B testing for applications such as web interface optimization, dynamic pricing, and targeted ad placement (see, bouneffouf2019survey for a survey of applications). In clinical trials, multi-stage or group sequential designs (wassmer2016group) have become standard, allowing early termination of experiments when strong evidence of positive or negative effects emerges. Economics has also seen a growing adoption of adaptive experimentation. For example, kasy2019adaptive and finan2022reinforcing develop new sequential experimentation strategies for use in development contexts. More recently, chapman2024dynamically introduced an optimal dynamic strategy for eliciting economic preferences.
@@
Despite these growing number of applications, determining optimal decision rules in adaptive experiments remains challenging due to the interactive nature of the data generating process. For instance, there is currently no well-established notion of optimal point estimation following an adaptive experiment, nor of inference procedures that remain valid at any point during the experiment. In this article, we address these challenges by deriving a continuous-time asymptotic representation for adaptive experiments. Specifically, we model the sequential data collection process using an empirical allocation process, which, at any given time $t$ specifies the fraction of time allocated to observing outcomes from a particular treatment. We then show that this empirical allocation process weakly converges to an allocation process in a limiting experiment, where signals consist of multiple Gaussian diffusions with unknown drifts, each corresponding to a different treatment arm. Then, by characterizing optimal decisions in the limit experiment, we can construct asymptotically optimal decision rules in the original experiment.
@@
The limit experiment greatly simplifies the characterization of optimal decisions due to two key properties. First, its state space is considerably smaller than that of the original experiment. In the limit experiment, the sufficient statistics are the past sample paths of the signal processes. However, in many applications, these statistics can be further reduced to just the current values of the signal and allocation processes, resulting in a state-space dimension of $2K$, where $K$ is the number of treatments. In contrast, the state space of the actual experiment encompasses all collected observations, making it substantially larger and more complex. Second, the limit experiment is far more tractable to analyze, as the properties of Gaussian diffusions in continuous time are well understood, allowing us to easily characterize optimal decisions in that setting.
@@
To illustrate the broad applicability of our framework, we use our representation theorem to derive optimal decision rules for several fundamental aspects of adaptive experimentation. In particular, we address: (1) the construction of optimal estimators following adaptive experiments, (2) the analysis of in-sample regret, and (3) the development of e-processes for anytime-valid inference. An e-process is a nonnegative supermartingale (under the null hypothesis) that provides a principled way to track statistical evidence against the null over the course of an experiment. We introduce the first definition of an e-process for multi-treatment adaptive experiments. This enables the design of algorithm-free anytime-valid tests, i.e., tests that maintain correct size even when the sampling strategy used in the adaptive experiment is unknown.
@@ 1.1. Related literature
In an important prior work, hirano2023asymptotic develop an asymptotic representation theorem for batched adaptive experiments, where sampling strategies are updated only a finite number of times over the course of the experiment. In this setting, they show that the policy rule and any statistic from the finite-sample experiment can be matched to a corresponding rule and statistic in a limit experiment involving Gaussian signals from each batch.
@@
Our asymptotic representation theorem is conceptually very different in terms of both the scope and the formulation. It applies to fully adaptive experiments and is expressed in terms of allocation processes rather than sequences of policy rules. This shift is essential, for in fully adaptive experiments, sequences of policy rules generally fail to admit weak limits. At the same time, our theory is more specialized in certain respects. It does not directly provide asymptotic representations for arbitrary statistics; instead, it characterizes the joint evolution of the score and the empirical allocation processes, which together fully determine the limit experiment. Obtaining representations for arbitrary statistics is more difficult in our setting because there is no straightforward coupling method for continuous-time processes that respects the required informational constraints (for example, ensuring that an anytime-valid test depends only on observed data).
@@
However, for many applications such comprehensive representation theorems are not always necessary. In practice, lower bounds on losses or risk can be established more directly using our representation theorem and standard change-of-measure arguments. For instance, in the case of point estimation, we show that the frequentist risk of any sequence of estimators is asymptotically bounded below by the risk of an estimator in the limit experiment that depends only on the terminal values of the signal and allocation processes. While this result does not establish a one-to-one mapping from finite-sample estimators to their counterparts in the limit experiment, such stronger representation is not essential for deriving lower bounds or constructing asymptotically optimal estimators.
@@
Beyond these technical contributions, the broader value of our approach is conceptual. This article provides the first general definition of adaptive experiments in continuous time through the lens of allocation processes. This formulation not only simplifies the analysis of adaptive experiments—continuous time being more tractable than discrete time—but also yields a sharp characterization of sufficient statistics. As discussed earlier, in many applications, including point estimation and anytime-valid inference, the sufficient statistics reduce to the current values of the signal and allocation processes, resulting in a finite-dimensional state-space representation. More generally, allocation processes offer a natural and flexible framework for representing adaptive experiments, and our formulation of e-processes is indeed most naturally articulated in terms of these processes.
@@
In terms of the style of asymptotic approximations, this article is most closely related to Le Cam’s (LeCam1979) work on stopping time representations. We extend his framework and build on his proof techniques to handle the additional complexities introduced by adaptive sampling.
@@
The systematic study of Gaussian diffusion approximations for adaptive experiments was initiated by fan2021diffusion and wager2021diffusion, who introduced diffusion asymptotics to analyze the behavior of adaptive algorithms such as Thompson Sampling. kalvit2021closer extended this approach to cover UCB algorithms, and adusumilli2025optimal further generalized the framework to characterize optimal bandit algorithms under both Bayesian and minimax regret criteria. adusumilli2025optimal also established that likelihood ratio processes from finite-sample adaptive experiments converge uniformly over time to their diffusion counterparts. While this convergence is algorithm-agnostic and well-suited for analyzing Bayesian decision criteria, it is insufficient for tasks such as anytime-valid inference, as it lacks a representation explicitly connecting the finite-sample algorithms to a suitable counterpart in the diffusion limit. This article fills that gap by providing precisely such a representation.
@@
Finally, this article also builds on earlier work by this author (adusumilli2023optimal) on optimal testing following adaptive experiments. The analysis of applications such as point estimation and anytime-valid inference draws on the strategies and proof techniques developed in that work.
@@ 2.2. General setup
Let $K$ denote the number of treatments under consideration. To simplify notation and proofs, we focus on the case $K=2$, though our results extend to any fixed $K$. The outcome under treatment $a\in\{0,1\}$ follows a parametric model $\{P_{\theta^{(a)}}^{(a)}\}$, where $\theta^{(a)}\in\mathbb{R}^{d}$ is an unknown parameter vector. For simplicity, we assume $\theta^{(0)}$ and $\theta^{(1)}$ have the same dimension, though this is not required for our results.
Since only one outcome is observed per unit, we can assume that outcomes are independent across treatments, conditional on $(\theta^{(1)},\theta^{(0)})$. Let $j=1,\dots,n$ index the experimental periods. We define time $t$ as the scaled number of periods, $t=j/n$, representing the fraction of the experiment completed.
The decision-maker (DM) employs a policy rule $\{\pi_{n,j}\}_{j}\equiv\{\pi_{n,\lfloor nt\rfloor}\}_{t}$, which prescribes the probability of assigning observation $j$ to treatment 1 based on past information. The treatment assignment follows $A_{j}\sim\text{Bernoulli}(\pi_{n,j})$. For the outcomes, it is useful to conceptualize a stack of potential observations $\bm{y}_{a}:=\{Y_{i}^{(a)}\}_{i=1}^{n}$ for each treatment, generated at the outset as i.i.d draws from $\{P_{\theta^{(a)}}^{(a)}\}$, but unobserved initially by the DM. This is the so-called stack-of-rewards model (lattimore2020bandit, Section 4.6). Each time a treatment is sampled, it can be imagined that the DM observes the top element of the corresponding treatment stack; this element is then taken out of consideration.
@@ 2.2.1. Empirical allocation processes
denote the fraction of (total) observations assigned to treatment $a$ up to time $t$. We term $\{q_{n,a}(\cdot)\}_{a}$ the empirical allocation process. One can interpret the policy rule $\{\pi_{n,j}\}_{j}$ as a function mapping the stacks $(\bm{y}_{1},\bm{y}_{0})$ and an exogenous random variable $U\sim\text{Uniform}[0,1]$ to the observed trajectory of $\{q_{n,a}(\cdot)\}_{a}$. The exogenous randomness accounts for probabilistic policy rules; specifically, we subsume the policy randomizations at all $n$ stages of the adaptive experiment into a single UU. [^1] Indeed, a single uniform random variable can be mapped into $n$ independent uniform random variables. The key informational constraint is that the event $\{q_{n,1}(t)\leq\gamma_{1},q_{n,0}(t)\leq\gamma_{0}\}$ depends only on the first $\lfloor n\gamma_{1}\rfloor,\lfloor n\gamma_{0}\rfloor$ observations from $(\bm{y}_{1},\bm{y}_{0})$ and the exogenous randomization $U$. Formally,
@@ 2.2.2. Local asymptotics
We are interested in the behavior of various statistics under local perturbations of the form $\theta_{0}^{(a)}+h^{(a)}/\sqrt{n}:h^{(a)}\in\mathbb{R}^{d}$}, where $\theta_{0}^{(a)}$ denotes a reference parameter. This focus is motivated by the fact that many decisions in adaptive experiments hinge on distinguishing between parameter values that are close to one another. For example, in the website optimization setting discussed earlier, bandit algorithms are employed precisely because the differences between website variants tend to be small. deng2013improving survey industry practices and report that typical differences in click-through rates are often around 1% or less. In such cases, bandit algorithms are employed for detecting subtle effects using as little data as possible. This naturally places us in the domain of local asymptotics.
@@ 2.2.3. Score processes
Knowledge of the the process, $z_{n,a}(\cdot)$, on the domain $[0,q]$ is equivalent to knowledge of the scores from the first $\left\lfloor nq\right\rfloor$ observations of the stack ${\bf y}_{a}$. We then define the score process for treatment $a$ as
@@ 2.3. The limit experiment
The primary result of this article establishes that any adaptive experiment is asymptotically Blackwell equivalent to a limit experiment characterized by Gaussian processes.
@@
The first two conditions on $q_{a}(\cdot)$ are straightforward. The third condition ensures that, aside from the exogenous randomization $U$, whether or not $q_{1}(t)\leq\gamma_{1}$ and $q_{0}(t)\leq\gamma_{0}$ hold can be determined based only on the sample paths, $\{z_{a}(s):s\leq\gamma_{a}\}$, of the cumulative past outcomes. It is the continuous time counterpart of the information constraint on empirical allocation processes, described in Section 2.2.1. The condition is analogous to the usual definition of a stopping time, but extended to a multi-dimensional setting.
Although the informational constraint is only explicitly required for $\gamma_{1}+\gamma_{0}\geq t$, the requirement $q_{1}(t)+q_{0}(t)=t$ almost surely implies that the event $\{q_{1}(t)\leq\gamma_{1},q_{0}(t)\leq\gamma_{0}\}$ is a measure-zero set when $\gamma_{1}+\gamma_{0}<t$. Consequently, because $\mathcal{G}_{\gamma_{1},\gamma_{0}}$ is an augmented filtration, this event is measurable with respect to $\mathcal{G}_{\gamma_{1},\gamma_{0}}$ for any $\gamma_{1}+\gamma_{0}<t$ as well.

### v2

@@ Abstract.
This article develops a continuous-time asymptotic framework for analyzing adaptive experiments—settings in which data collection and treatment assignment evolve dynamically in response to incoming information. A key challenge in analyzing fully adaptive experiments, where the assignment policy is updated after each observation, is that the sequence of policy rules often lack a well-defined asymptotic limit. To address this, we focus instead on the empirical allocation process, which captures the (normalized) number of observations assigned to each treatment over time. We show that, under general conditions, any adaptive experiment and its associated empirical allocation process can be approximated by a limit experiment defined by Gaussian diffusions with unknown drifts and a corresponding continuous-time allocation process. This limit representation facilitates the analysis of optimal decision rules by reducing the dimensionality of the state-space and exploiting the tractability of Gaussian diffusions. We apply the framework to derive optimal estimators, analyze in-sample regret for adaptive experiments, and construct e-processes for anytime-valid inference. Notably, we introduce the first definition of any-time and any-experiment valid inference for multi-treatment settings.
This version:
@@ 1. Introduction
Adaptive experiments are experiments where data is collected and analyzed continuously, allowing for adjustments or decisions to be made on an ongoing basis. Originating from early work by Wald (1947), Arrow et al. (1949), and others, adaptive experimentation has evolved to encompass a wide range of strategies, including bandit experiments, A/B testing, costly sampling, and best-arm identification. These strategies are now widely used across various fields. For instance, tech companies frequently employ bandit algorithms and A/B testing for applications such as web interface optimization, dynamic pricing, and targeted ad placement (see, Bouneffouf and Rish, 2019 for a survey of applications). In clinical trials, multi-stage or group sequential designs (Wassmer and Brannath, 2016) have become standard, allowing early termination of experiments when strong evidence of positive or negative effects emerges. Economics has also seen a growing adoption of adaptive experimentation. For example, Kasy and Sautmann (2021) and Finan and Pouzo (2022) develop new sequential experimentation strategies for use in development contexts. More recently, Chapman et al. (2024) introduced an optimal dynamic strategy for eliciting economic preferences.
@@
Despite these growing number of applications, determining optimal decision rules in adaptive experiments remains challenging due to the interactive nature of the data generating process. For instance, there is currently no well-established notion of optimal point estimation following an adaptive experiment, nor of inference procedures that remain valid at any point during the experiment. In this article, we address these challenges by deriving a continuous-time asymptotic representation for adaptive experiments. Specifically, we model the sequential data collection process using an empirical allocation process, which, at any given time $t$ specifies the fraction of time allocated to observing outcomes from a particular treatment. We then show that this empirical allocation process weakly converges to an allocation process in a limiting experiment, where signals consist of multiple Gaussian processes with unknown drifts, each corresponding to a different treatment arm. Then, by characterizing optimal decisions in the limit experiment, we can construct asymptotically optimal decision rules in the original experiment.
@@
The limit experiment greatly simplifies the characterization of optimal decisions due to two key properties. First, its state space is considerably smaller than that of the original experiment. In the limit experiment, the sufficient statistics are the past sample paths of the signal processes. However, in many applications, we show that these statistics can be further reduced to just the current values of the signal and allocation processes, resulting in a state-space dimension of $2K$, where $K$ is the number of treatments. In contrast, the state space of the actual experiment encompasses all collected observations, making it substantially larger and more complex. Second, the limit experiment is far more tractable to analyze, as the properties of Gaussian diffusions in continuous time are well understood, allowing us to easily characterize optimal decisions in that setting.
@@
To illustrate the broad applicability of our framework, we use our representation theorem to derive optimal decision rules for several fundamental aspects of adaptive experimentation. In particular, we address: (1) the construction of optimal estimators following adaptive experiments, (2) the analysis of in-sample regret, and (3) the development of e-processes for anytime-valid inference. An e-process is a nonnegative supermartingale (under the null hypothesis) that continuously tracks statistical evidence against the null over the course of an experiment. We introduce the first definition of an e-process for multi-treatment adaptive experiments. This enables the design of algorithm-free anytime-valid tests, i.e., tests that maintain correct size even when the sampling strategy used in the adaptive experiment is unknown.
@@ 1.1. Related literature
In an important prior work, Hirano and Porter (2023) develop an asymptotic representation theorem for batched adaptive experiments, where sampling strategies are updated only a finite number of times over the course of the experiment. In this setting, they show that the policy rule and any statistic from the finite-sample experiment can be matched to a corresponding rule and statistic in a limit experiment involving Gaussian signals from each batch.
@@
Our asymptotic representation theorem is very different in terms of both the scope and the formulation. It applies to fully adaptive experiments and is expressed in terms of allocation processes rather than sequences of policy rules. This shift is essential, for in fully adaptive experiments, sequences of policy rules generally fail to admit weak limits. At the same time, our theory is more specialized in certain respects. It does not directly provide asymptotic representations for arbitrary statistics; instead, it characterizes the joint evolution of the score and the empirical allocation processes, which together fully determine the limit experiment. Obtaining representations for arbitrary statistics is more difficult in our setting because there is no straightforward coupling method for continuous-time processes that respects the required informational constraints (for example, ensuring that an anytime-valid test depends only on observed data).
@@
However, for many applications such comprehensive representation theorems are not always necessary. In practice, lower bounds on losses or risk can be established using our representation theorem and standard change-of-measure arguments. For instance, in the case of point estimation, we show that the frequentist risk of any sequence of estimators is asymptotically bounded below by the risk of an estimator in the limit experiment that depends only on the terminal values of the signal and allocation processes. While this result does not establish a one-to-one mapping from finite-sample estimators to their counterparts in the limit experiment, such stronger representation is not essential for deriving lower bounds or constructing asymptotically optimal estimators.
@@
Beyond these technical contributions, the broader value of our approach is conceptual. This article provides the first general definition of adaptive experiments in continuous time through the lens of allocation processes. This formulation not only simplifies the analysis of adaptive experiments—continuous time being more tractable than discrete time—but also yields a sharp characterization of sufficient statistics. As discussed earlier, in many applications, including point estimation and anytime-valid inference, the sufficient statistics reduce to the current values of the signal and allocation processes, resulting in a finite-dimensional state-space. More generally, allocation processes offer a natural and flexible framework for representing adaptive experiments, and our formulation of e-processes is indeed most naturally articulated in terms of these processes.
@@
In terms of the style of asymptotic approximations, this article is most closely related to Le Cam’s (1979) work on stopping time representations. We extend his framework and build on his proof techniques to handle the additional complexities introduced by adaptive sampling.
@@
The systematic study of Gaussian process approximations for adaptive experiments was initiated by Fan and Glynn (2021) and Kuang and Wager (2024), who introduced diffusion asymptotics to analyze the behavior of adaptive algorithms such as Thompson Sampling. Kalvit and Zeevi (2021) extended this approach to cover UCB algorithms, and Adusumilli (2025b) further generalized the framework to characterize optimal bandit algorithms under both Bayesian and minimax regret criteria. Adusumilli (2025b) also established that likelihood ratio processes from finite-sample adaptive experiments converge uniformly to their diffusion counterparts. While this convergence is algorithm-agnostic and well-suited for analyzing Bayesian decision criteria, it is insufficient for tasks such as anytime-valid inference, as it lacks a representation explicitly connecting the finite-sample algorithms to a suitable counterpart in the diffusion limit. This article fills that gap by providing precisely such a representation.
@@
Finally, this article also builds on earlier work by this author (Adusumilli, 2023) on optimal testing following adaptive experiments. The analysis of applications such as point estimation and anytime-valid inference draws on the strategies and proof techniques developed in that work.
@@ 2.2. General setup
Let $K$ denote the number of treatments under consideration. To simplify notation and proofs, we focus on the case $K=2$, though our results extend to any fixed $K$. The outcome $Y^{(a)}$ under treatment $a\in\{0,1\}$ follows a parametric model $\{P_{\theta^{(a)}}^{(a)}\}$, where $\theta^{(a)}\in\mathbb{R}^{d}$ is an unknown parameter. For simplicity, we assume $\theta^{(0)}$ and $\theta^{(1)}$ have the same dimension, though this is not required for our results.
Since only ever observe one potential outcome $Y^{(a)}$ per unit, we can assume $\{P_{\theta^{(a)}}^{(a)}\}_{a}$ are independent across $a$, conditional on $(\theta^{(1)},\theta^{(0)})$. Let $j=1,\dots,n$ index the experimental periods. We define time $t$ as the scaled number of periods, $t=j/n$, representing the fraction of the experiment completed.
The decision-maker (DM) employs a policy rule $\{\pi_{n,j}\}_{j}\equiv\{\pi_{n,\lfloor nt\rfloor}\}_{t}$, which prescribes the probability of assigning observation $j$ to treatment 1 based on past information. The treatment assignment follows $A_{j}\sim\text{Bernoulli}(\pi_{n,j})$. For the outcomes, it is useful to conceptualize a stack of observations $\bm{y}_{a}:=\{Y_{i}^{(a)}\}_{i=1}^{n}$ for each treatment, generated at the outset as i.i.d draws from $\{P_{\theta^{(a)}}^{(a)}\}$, but unobserved initially by the DM. This is the so-called stack-of-rewards model (Lattimore and Szepesvári, 2020, Section 4.6). Each time a treatment is sampled, it can be imagined that the DM observes the top element of the corresponding treatment stack; this element is then taken out of consideration.
@@ 2.2.1. Empirical allocation processes
denote the fraction of (total) observations assigned to treatment $a$ up to time $t$. We term $\{q_{n,a}(\cdot)\}_{a}$ the empirical allocation process. One can interpret the policy rule $\{\pi_{n,j}\}_{j}$ as a function mapping the stacks $(\bm{y}_{1},\bm{y}_{0})$ and an exogenous random variable $U\sim\text{Uniform}[0,1]$ to the observed trajectory of $\{q_{n,a}(\cdot)\}_{a}$. The exogenous randomness accounts for probabilistic policy rules; specifically, we subsume the policy randomizations at all $n$ stages of the adaptive experiment into a single UU. [^1] Indeed, a single uniform random variable can be mapped into countably infinite independent uniform random variables. The key informational constraint is that the event $\{q_{n,1}(t)\leq\gamma_{1},q_{n,0}(t)\leq\gamma_{0}\}$ depends only on the first $\lfloor n\gamma_{1}\rfloor,\lfloor n\gamma_{0}\rfloor$ observations from $(\bm{y}_{1},\bm{y}_{0})$ and the exogenous randomization $U$. Formally,
@@ 2.2.2. Local asymptotics
We are interested in the behavior of various statistics under local perturbations of the form $\theta_{0}^{(a)}+h^{(a)}/\sqrt{n}:h^{(a)}\in\mathbb{R}^{d}$}, where $\theta_{0}^{(a)}$ denotes a reference parameter. This focus is motivated by the fact that many decisions in adaptive experiments hinge on distinguishing between parameter values that are close to one another. For example, in the website optimization setting discussed earlier, bandit algorithms are employed precisely because the differences between website variants tend to be small. Deng et al. (2013) survey industry practices and report that typical differences in click-through rates are often around 1% or less. In such cases, bandit algorithms are employed for detecting subtle effects using as little data as possible. This naturally places us in the domain of local asymptotics.
@@ 2.2.3. Score processes
Knowledge of the process, $z_{n,a}(\cdot)$, on the domain $[0,q]$ is equivalent to knowledge of the scores from the first $\left\lfloor nq\right\rfloor$ observations of the stack ${\bf y}_{a}$. We then define the score process for treatment $a$ as
@@ 2.3. The limit experiment
The primary result of this article establishes that any adaptive experiment is asymptotically Blackwell equivalent to a limit experiment characterized by Gaussian diffusions.
@@
Definition 1 addresses the technical challenge of defining adaptive experiments in continuous time without recourse to policy rules, which are typically non-measurable (see Section 2.3.2 below). Instead, it defines allocation processes abstractly through their functional properties. The first two conditions on $q_{a}(\cdot)$ are straightforward. The third condition ensures that, aside from the exogenous randomization $U$, whether or not $q_{1}(t)\leq\gamma_{1}$ and $q_{0}(t)\leq\gamma_{0}$ hold can be determined based only on the sample paths, $\{z_{a}(s):s\leq\gamma_{a}\}$, of the signal processes. [^3] Although the informational constraint is only explicitly required for $\gamma_{1}+\gamma_{0}\geq t$, the requirement $q_{1}(t)+q_{0}(t)=t$ almost surely implies that the event $\{q_{1}(t)\leq\gamma_{1},q_{0}(t)\leq\gamma_{0}\}$ is a measure-zero set when $\gamma_{1}+\gamma_{0}<t$. Consequently, because $\mathcal{G}_{\gamma_{1},\gamma_{0}}$ is an augmented filtration, this event is measurable with respect to $\mathcal{G}_{\gamma_{1},\gamma_{0}}$ for any $\gamma_{1}+\gamma_{0}<t$ as well. It is the continuous time counterpart of the information constraint on empirical allocation processes, described in Section 2.2.1. The condition is analogous to the usual definition of a stopping time, but extended to a multi-dimensional setting.